# Toward Evaluation Frameworks for Multi-Agent Scientific AI Systems

### 18 Mar 2026 v1

@@ I Introduction
We consider a class of scientific AI systems that integrate large language models with external tools, simulation capabilities, and iterative reasoning. Such systems have been proposed in various forms in recent literature and industry prototypes [1, 12, 18, 11]. Here, we will focus on workflows supporting (or augmenting) the work of physicists in such areas as condensed matter physics and quantum information.
@@
In this report, we attempt to systematize this problem. Our goal is not to claim that there exists a perfect way to measure scientific capabilities; not even that there exists a single metric that can capture those; but rather to explore how practical tests can be designed to measure some of the non-trivial aspects of a scientific AI system. These include, among others, the ability to generate novel hypotheses, detect mistakes in existing arguments, synthesize information from multiple sources, and reason about partially specified or unfamiliar physical scenarios.
@@
In the following sections, we discuss the conceptual challenges of benchmarking scientific (multi-)agentic systems and propose a set of evaluation strategies to probe selected aspects of those systems. The emphasis throughout the work is on designing tests that go beyond simple retrieval or computation and instead assess aspects of reasoning central to the practice of science.
II The Benchmarking Problem
@@
III Benchmark Design Principles
@@ IV.1 Replication Benchmarks
IV.2 Error Detection Benchmarks
@@
IV.3 Scientific Reasoning Benchmarks
IV.3.1 Future Research Directions
@@ IV.3.2 Assumption Enumeration
IV.3.3 Minimal Counterexample Construction
@@ IV.3.4 Technique Recommendation
IV.4 Discovery-Style Benchmarks
IV.4.1 Fabricated Phenomena
@@
IV.4.2 Solving Novel Problems
@@
V Benchmark Construction Strategies
@@
V.1 Planted-Structure Problems
@@
V.2 Constraint-Modified Problems
@@
V.3 Dynamic Difficulty Families
@@
V.4 Self-Consistency Tests
@@
V.5 Multi-Turn Scientific Dialogue
@@
Example: “I have a bipartite quantum system.” (System should ask about dimensions, state preparation, available measurements.) “It is a two-qubit system.” (System should identify the relevant framework.) “I measured these Bell correlations: [data].” (System should analyze whether they violate a Bell inequality.) “Now I add a third qubit.” (System should adapt to the multipartite setting.) We might continue until the system has fully characterized a planted scenario.
This tests the system’s ability to ask the right questions (not just answer given ones), mirrors real scientific practice. It is also related with the notion of “understanding” of what the system does not know.
@@
VI System Evaluation Protocol
VI.1 Ablation Studies
We should test how the system behaves when we disable specific components, e.g., web search, knowledge graph retrieval, code sandbox. This ablation structure directly addresses the retrieval–reasoning confound, e.g., if performance drops only slightly when we disable retrieval, the benchmark is genuinely testing reasoning. Contrary, if it drops to near-zero, the benchmark is mostly testing retrieval.
This will allow us to collect diagnostic information about the benchmark itself (to learn what what a given benchmark tests).
VI.2 Tool-Use Analysis
For any benchmark problem P where a known algorithmic solution exists, the system can pursue two strategies: (A) retrieve the algorithm and implement it, or (B) reason from first principles. Strategy A tests retrieval plus code generation; Strategy B tests multi-step scientific reasoning. A well-designed agentic system should use retrieval when a known algorithm exists and construct solution from first principle only if such approach is needed. By checking the reasoning trace, we can verify this.
VI.3 Process-Level Evaluation
Checking final answer is not enough. The way how the system solves problems matter. We must capture the trace of the solution and examine, whether the systems invokes the right definitions, follows a valid logical chain, produce rigorous arguments, etc.
VI.4 Scaling Curves
@@
VI.5 Point Tests
We must develop a collection of “unit tests” for individual subsystems. These are essential for diagnosing potential failures. We should test both tool selection and performance of specific tools such as symbolic regression, proof tooling, etc.
VI.6 Integration Tests
We must test whether components work together. There might be many possible combinations (literature search $\rightarrow$ symbolic regression $\rightarrow$ proof tooling, or literature search $\rightarrow$ numerical simulation $\rightarrow$ synthesis). Testing short, well defined pipelines can give us more control (and more precise-localized feedback) than running end-to-end tasks where we don’t know how different elements of the system contributs to the overall sucess (or failur).
VI.7 Red Teaming and Adversarial Robustness Tests
We might create a tailored, challenging tests to expose particular (expected) shortcomings of the system. For example, we might ask a question that is similar to (but at its essence, very different from) a known popular question. We can test whether the system will recognize the difference or whether it will provide answer to the popular question (an example of “mode collapse” in LLMs).
This idea is close to the concept of Red Teaming, where the focus would be on creating test examples that aim to exploits or break the system (e.g., by supplying specific strings of numbers to cause spontaneous broad misalignment as described in [2]).
VI.8 Expert-Sourced Problems
We might attempt to create similar collection of problems, based on never published results, similar to how it was done in [19].
VII Human-Centered Evaluation
@@
To get some initial intuition, we conducted several free-form interviews with professionals ranging from quantum researchers to quantum software engineers to understand their experiences with AI systems, their current pain points, and their needs. The questions covered their workplace, typical tasks, which tasks they consider essential versus delegatable, how they currently work with AI, and what role they would like AI to play.
VII.1 Selected Quotes
@@ VII.3 Sentiment Analysis
VII.4 Implications for Benchmark Design
@@ VIII Feasibility Study
While the pipeline still requires improvement (we can add a verification loop, we can match future directions with multiple categories, we can filter out papers that were already discussed elsewhere, etc.), this simple test shows that building such a dynamic database of concepts is feasible.
This direction might also be similar to [6].
IX Conclusion

### 05 Apr 2026 v2

@@ I Introduction
We consider a class of scientific AI systems that integrate large language models with external tools, simulation capabilities, and iterative reasoning. Such systems have been proposed in various forms in recent literature and industry prototypes [3, 21, 36, 20]. Here, we will focus on a general class of such systems, supporting (or augmenting) the work of physicists in such areas as condensed matter physics and quantum information. An architectural overview of such a system might be similar to the diagram depicted in Fig. 1.
Figure 1: An architectural overview of a generalized scientific AI system for physics research assistance. The interaction layer may take multiple forms, including a conversational assistant, a research workspace, an augmented text/code editor, a background agent, or a different form. The orchestration layer coordinates actions, routes problems, maintains state, plans workflows, executes tools, etc. The data layer is not restricted to offline databases; we assume that our system can and will perform a life search on the internet. The validation layer is not executed only at the end, it might be in constant dialog with the orchestration layer providing feedback and critic of particular steps.
@@
In this report, we attempt to systematize this problem. Our goal is not to claim that there exists a perfect way to measure scientific capabilities of an automatic (or semi-automatic) artificial systems; not even that there exists a single metric that can capture that concept; but rather to explore how practical tests can be designed to measure some of the non-trivial aspects of scientific AI systems. These aspects include, among others, the ability to generate novel hypotheses, detect mistakes in existing arguments, synthesize information from multiple sources, and reason about partially specified or unfamiliar physical scenarios.
II Existing Works
Early efforts focused on testing LLMs on sets of closed-form questions, often multiple-choice, drawn from standardised examinations. MMLU [13] introduced a broad 57-subject evaluation covering undergraduate and professional topics, while MMLU-Pro tightened the task by providing ten answer choices and selecting questions that require multi-step reasoning rather than simple recall. GPQA [25] pushed further toward expert difficulty: the 448 graduate-level questions in biology, physics, and chemistry are intentionally “Google-proof,” designed so that non-expert PhD holders who may search the web freely still score only around 34%, while domain specialists reach 65–74%. The Diamond subset retains only those problems that experts answered correctly, but generalists failed, which made the set widely popular. Specialized physics benchmarks have followed a similar trajectory: UGPhysics [34] targets undergraduate-level problems across many sub-disciplines; PhysReason [38] extends evaluation to multi-step reasoning chains with step-level scoring; SciBench [32] focuses on college-level quantitative problem solving; and OlympiadBench [12] covers competition-level problems across multiple scientific fields.
These benchmarks share a common structure and a common limitation. The closed-form, single-shot format privileges knowledge recall and procedural computation over the open-ended, iterative reasoning that characterizes real scientific practice. Performance on such tests is, moreover, approaching saturation: frontier models now exceed PhD-level accuracy on GPQA Diamond (reaching above 90% on some evaluations), which both demonstrates progress and signals that harder, more realistic tests are needed. Closed-form benchmarks also encourage models to make guesses (hallucinate answers), engage in reward hacking, and, overall, often do not distinguish between retrieval with actual reasoning.
A newer generation of benchmarks attempts to evaluate richer scientific capabilities beyond question answering. CURIE [9] tests long-context scientific understanding, drawing on full research papers in materials science, condensed matter physics, quantum computing, and related fields to probe information extraction, concept tracking, and cross-domain synthesis. The benchmark requires models to reason over complete papers rather than isolated snippets. FrontierScience [31] targets PhD-level open-ended research problems in physics, chemistry, and biology. Problems are specified with structured rubrics that decompose expected answers into individually gradable components. CritPt [40] adopts a complementary approach: it focuses specifically on frontier physics research, using unpublished problems sourced from domain experts to circumvent retrieval contamination. While that benchmark may reflect the current state-of-the-art approach, it is costly to generate, subject to individual researchers’ biases, and has an opaque evaluation procedure.
ResearcherBench [33] targets deep AI research systems that autonomously search the web, synthesize multi-source evidence, and produce structured research reports on frontier questions. The problem of data contamination is discussed in [40], and the idea of generating "interesting scientific ideas" using knowledge graphs and LLMs is presented in [11].
Others explored the idea of scientific paper replication. PaperBench [29] evaluates AI agents on 20 ICML 2024 papers, requiring them to understand each paper’s contributions, write a codebase from scratch, and execute experiments successfully. The task is decomposed into 8,316 individually gradable sub-tasks using rubrics co-developed with the original authors. The best-tested agent at launch (Claude 3.5 Sonnet with open-source scaffolding) achieved a replication score of 21.0%, compared to 41.4% for a human ML PhD baseline given 48 hours. While PaperBench focuses on ML research, it establishes the rubric-graded, agentic-scaffold evaluation methodology that is broadly applicable.
ReplicationBench [37] applies a similar philosophy to astrophysics, constructing paper-scale expert-validated tasks that measure both faithfulness (adherence to the original methodology) and correctness (technical accuracy). Frontier models score below 20% even on this domain-specific benchmark, and analysis of failure trajectories reveals a rich taxonomy of error modes. Axiomatic AI’s Lemma product [5] takes a commercial approach to the same idea, automatically verifying mathematical claims extracted from papers against formal symbolic computation.
The formalization of mathematical reasoning in proof assistants such as Lean has created a family of benchmarks with provably correct evaluation: either the proof compiles, or it does not. AlphaProof [15], combining reinforcement learning with Lean’s proof checker, reached silver-medal performance on IMO 2024 problems, and subsequent systems, including Harmonic’s Aristotle and DeepSeek-Prover [1], have since reached gold-medal levels on IMO 2025. The miniF2F benchmark [39], which collects Olympiad-level problems in a formal language, now serves as a standard measure of automated theorem-proving capability. More recently, agentic frameworks such as APOLLO [23] have combined LLM generation with symbolic repair loops to achieve over 84% on miniF2F with small models.
This line of work is highly relevant but also meaningfully different from the physics-focused benchmarks we discuss. Formal mathematics operates in closed, axiomatic systems where every inference step can be verified mechanically. Theoretical physics, by contrast, requires connecting formal reasoning to physical intuition, experiment, and domain knowledge that cannot easily be encoded in a proof assistant. An AI system that excels at formal proof search may still struggle to select the right physical model, identify the relevant regime, or interpret the significance of a numerical result. We therefore treat formal mathematical reasoning as an important component of the system under test, particularly via the proof tooling and CAS subsystems, rather than as a sufficient benchmark in itself.
The AI Scientist [19] introduced a complete pipeline for autonomous scientific discovery in machine learning: generating hypotheses, writing and executing experimental code, and producing a full paper with automated review. Its successor, AI Scientist-v2 [35], removed the requirement for human-authored code templates and introduced an agentic tree-search strategy, producing the first entirely AI-generated workshop paper to pass peer review. These systems demonstrate that end-to-end autonomous research is no longer purely speculative, at least within the constrained setting of ML experiments on standard benchmarks.
At the industry level, all major AI labs have announced efforts in this direction. Google DeepMind released an “AI Co-Scientist” system [10] designed to assist with hypothesis generation and experimental design. OpenAI established an “OpenAI for Science” initiative [22] and has highlighted GPT-5’s performance on science-specific benchmarks. Anthropic launched Claude for Life Sciences [2], integrating Claude with tools such as PubMed, Benchling, and 10x Genomics to support the full research pipeline from literature review to regulatory submission.
@@ II Existing Works
Despite the rapid pace of this work, a gap remains between what these systems demonstrate and what practicing scientists actually need. In the following sections, we discuss the conceptual challenges of benchmarking scientific (multi-)agentic systems and propose a set of evaluation strategies to probe selected aspects of those systems. The emphasis throughout the work is on designing tests that go beyond simple retrieval or computation and instead assess aspects of reasoning central to the practice of science.
III The Benchmarking Problem
@@
IV Benchmark Design Principles
@@ V.1 Replication Benchmarks
V.2 Error Detection Benchmarks
@@
V.3 Scientific Reasoning Benchmarks
V.3.1 Future Research Directions
@@ V.3.2 Assumption Enumeration
V.3.3 Minimal Counterexample Construction
@@ V.3.4 Technique Recommendation
V.4 Discovery-Style Benchmarks
V.4.1 Fabricated Phenomena
@@
V.4.2 Solving Novel Problems
@@
VI Benchmark Construction Strategies
@@
VI.1 Planted-Structure Problems
@@
VI.2 Constraint-Modified Problems
@@
VI.3 Dynamic Difficulty Families
@@
VI.4 Self-Consistency Tests
@@
VI.5 Multi-Turn Scientific Dialogue
@@
Example: “I have a bipartite quantum system.” (The system should ask about dimensions, state preparation, and available measurements.) “It is a two-qubit system.” (The system should identify the relevant framework.) “I measured these Bell correlations: [data].” (The system should analyze whether they violate a Bell inequality.) “Now I add a third qubit.” (The system should adapt to the multipartite setting.) We might continue until the system has fully characterized a planted scenario.
This tests the system’s ability to ask the right questions (not just answer given ones), mirrors real scientific practice. It is also related to the notion of “understanding” of what the system does not know.
@@
VII System Evaluation Protocol
A multi-agent system is inherently complex and requires a multi-faced approach for testing. Some of the approaches should include:
Ablation Studies.
We should test how the system behaves when we disable specific components, e.g., web search, knowledge graph retrieval, code sandbox. This ablation structure directly addresses the retrieval–reasoning confound, e.g., if performance drops only slightly when we disable retrieval, the benchmark is genuinely testing reasoning. Contrary, if it drops to near-zero, the benchmark is mostly testing retrieval. This will allow us to collect diagnostic information about the benchmark itself (to learn what a given benchmark tests).
Tool-Use Analysis.
For any benchmark problem P where a known algorithmic solution exists, the system can pursue two strategies: (A) retrieve the algorithm and implement it, or (B) reason from first principles. Strategy A tests retrieval plus code generation; Strategy B tests multi-step scientific reasoning. A well-designed agentic system should use retrieval when a known algorithm exists and construct a solution from first principles only if such an approach is needed. By checking the reasoning trace, we can verify this.
Process-Level Evaluation.
Checking the final answer is not enough. The way emphhow the system solves problems matters. We must capture the trace of the solution and examine whether the system invokes the right definitions, follows a valid logical chain, produces rigorous arguments, etc.
Scaling Curves.
@@ VII System Evaluation Protocol
Point Tests.
We must develop a collection of “unit tests” for individual subsystems. These are essential for diagnosing potential failures. We should test both tool selection and performance of specific tools, such as symbolic regression, proof tooling, etc.
Integration Tests.
We must test whether components work together. There might be many possible combinations (literature search $\rightarrow$ symbolic regression $\rightarrow$ proof tooling, or literature search $\rightarrow$ numerical simulation $\rightarrow$ synthesis). Testing short, well-defined pipelines can give us more control (and more precise, localized feedback) than running end-to-end tasks where we don’t know how different elements of the system contribute to the overall success (or failure).
Red Teaming and Adversarial Robustness Tests.
We might create tailored, challenging tests to expose particular (expected) shortcomings of the system. For example, we might ask a question that is similar to (but at its essence, very different from) a known popular question. We can test whether the system will recognize the difference or whether it will provide an answer to the popular question (an example of “mode collapse” in LLMs). This idea is close to the concept of Red Teaming, where the focus would be on creating test examples that aim to exploit or break the system (e.g., by supplying specific strings of numbers to cause spontaneous broad misalignment as described in [4]).
Expert-Sourced Problems.
We might attempt to create a similar collection of problems, based on never-published results, similar to how it was done in [40].
VIII Human-Centered Evaluation
@@
To gain initial intuition, we conducted several open-ended interviews with professionals ranging from quantum researchers to quantum software engineers to understand their experiences with AI systems, current pain points, and needs. The questions covered their workplace, typical tasks, which tasks they consider essential versus delegatable, how they currently work with AI, and what role they would like AI to play.
VIII.1 Selected Quotes
@@ VIII.3 Sentiment Analysis
VIII.4 Implications for Benchmark Design
@@ IX Feasibility Study
While the pipeline still requires improvement (e.g., adding a verification loop, aligning future directions with multiple categories, filtering out papers that have already been discussed elsewhere, etc.), this simple test shows that building such a dynamic database of concepts is feasible (cf. Ref. [11]).
X Conclusion